# Exploring Semi-Supervised Learning for Online Mapping

### v1

@@ III-F Teacher Fusion
To maximize overlap between additional samples and the area of the main sample, and to ensure that these additional samples provide new information, we select them from positions at randomly sampled distances relative to the current main frame. This approach mitigates issues associated with time-based sample selection, such as the ego vehicle being stationary—which offers no additional information—or moving at very high speeds, resulting in little overlap.
IV Experiments
To maximize the effectiveness of adding unlabelled data in the online mapping setting, we optimize parameters and design choices for each of the discussed methods and assess the relative impact of different components through ablation studies. Additionally, we examine the usefulness of our method by training on one set of cities and evaluating on another, assessing its capability for domain adaptation.
IV-A Experimental Setup
We employ a ResNet50 [20] as image encoder $f^{\text{enc}}_{\theta}$, LSS [36] for lifting features to BEV $f^{\text{lift}}_{\theta}$, and a lightweight convolutional bottleneck decoder inspired by [18] as $f^{\text{dec}}_{\theta}$. We perform multi-label prediction for pedestrian crossings, lane dividers, and road boundaries and use Focal Loss [28] as the supervised loss $\ell_{\text{sup}}$ on the labelled dataset and $\ell_{\text{cls}}$ for the unlabelled data. The region of interest follows [30] spanning ±$\pm 45$m longitudinally and ±$\pm 15$m laterally. Each cell has a side length of $0.3$m. For brevity, we report the mean Intersection over Union (mIoU) across these three classes. To account for the stochasticity in the results from individual runs, we either report the mean and standard deviation of the mIoU for three runs with the same hyperparameters or multiple nearby values of the hyperparameter. Empirically, we note that our setup is robust to a range of unlabelled weights $\omega_{\text{cls}}$ spanning $0.25$ to $5.0$. We choose to follow [14] and use $\omega_{\text{cls}}=1$. We further follow [21] and linearly ramp-up $\omega_{\text{cls}}$, and when applicable $\omega_{\text{feat}}$, over the first third of the training. Unless stated otherwise, no thresholding or sharpening is used in the experiments. The models, unless stated otherwise, are trained and evaluated using the Argoverse 2 (AV2) dataset [45] with the geographically disjoint splits proposed by [27] using surround-view camera images as input and use the provided HD maps as labels. We generally experiment using two utilisation rates of labelled data, namely $2.5\%$ and $10\%$, corresponding to $18$ and $70$ labelled sequences respectively.
IV-B Augmentations
TABLE I: Performance (mIoU) of various augmentations.
Augmentation Label utilisation
@@ IV-B Augmentations
$21.9^{\pm 1.0}$ $31.9^{\pm 0.1}$
$24.4^{\pm 1.2}$ $33.9^{\pm 0.2}$
$25.5^{\pm 0.7}$ $34.3^{\pm 0.1}$
Empirically we observe that using a high probability of applying the augmentation yields better performance and choose always to apply the strong augmentations to the student. For the CutOut augmentation, we find that removing $25\%$ of the image yields good results and is thus used for all experiments. For BEV feature dropout (BEVDrop) we follow [21, 48] and drop $50\%$ of the features. The results in Tab. I demonstrate that increasing the difficulty of the task for the student model is crucial also for online mapping. With $10\%$ label utilisation, Photometric augmentation alone is sufficient for significant model improvement. However, with $2.5\%$ label utilisation, additional augmentations are required. For both label utilisation ratios, applying Photometric, CutOut, and BEVDrop yields the best performance, and we adopt this configuration as the Baseline in further experiments. Specifically, we observe a significant performance improvement, with mIoU increasing from $12.1$ to $25.5$ for the $2.5\%$ label utilisation and $21.3$ to $34.3$ for the $10\%$ label utilization. Furthermore, contrary to the findings in [21], we find that removing parts of the images, i.e. CutOut, yields better performance than dropping an entire camera and masking its field of view in the loss, as done with CamDrop.
IV-C Sharpening and Thresholding
Fig. 2 illustrates the effects of sharpening and thresholding. As shown on the left, supervising with only the most confident pseudo-labels does not yield any additional performance boost. Using a low threshold can marginally improve performance, and we opt to use $0.6$ as the threshold in further experiments. Tab. II displays the results over multiple runs using hard and soft labels. We note that there is no significant change in performance using sharpening, which differs from other domains where sharpening has shown to be very effective, usually with high thresholds [37, 48, 44]. While applying sharpening with a temperature parameter can improve performance, right column in Fig. 2, the gains are marginal and introduce an additional hyperparameter.
Figure 2: Thresholding and sharpening utilising $2.5\%$ and $10\%$ of the labels. Baseline refers to using only strong augmentations.
TABLE II: mIoU for sharpening methods using threshold $0.6$.
Technique Label utilisation
@@ IV-C Sharpening and Thresholding
$25.4^{\pm 0.8}$ $34.6^{\pm 0.1}$
IV-D Feature Consistency
@@
We compare the performance using Mean Squared Error (MSE) and Cosine Similarity (COS) losses with $\omega_{\text{cls}}=1$ and varying feature loss weight $\omega_{\text{feat}}$. In Fig. 3 the results are displayed using features from before (Pre) or after (Post) the BEV decoder $f^{\text{dec}}_{\theta}$ as inputs to $\ell_{\text{feat}}$. When utilizing $2.5\%$ of the labels the results are noisy, and no clear benefit from the additional supervision is observed across the setups. However, for the $10\%$ label utilisation, COS consistency provides a slight performance gain and demonstrates greater stability than the MSE loss. While prior BEV segmentation works opted to use MSE [55, 14], we believe COS to be the better choice.
Figure 3: Feature Similarity using Mean Square Error and Cosine Similarity with varying weights. Baseline refers to using only strong augmentations. Pre and Post refers to using $\mathbf{z}_{\text{bev}}$ or $\mathbf{\tilde{z}}_{\text{bev}}$ respectively.
IV-E Teacher Fusion
@@
To enhance pseudo-labels, we fuse teacher predictions from multiple samples. We use two additional samples, as we empirically found that adding more samples does not provide further benefits. Therefore, we focus on examining the influence of the maximum distance to these additional samples. The results in Fig. 4 show that even our simple fusion of predicted probabilities generally improves performance, both with and without thresholding. The fusion of BEV features, on the other hand, shows no significant benefit. For $2.5\%$ label utilisation we observe that increasing the range to above $20$m is required to improve performance, whereas for $10\%$ the maximum range is not as important. In further experiments, we will use $30$m as it works well for both utilisation ratios. While the simple probability fusion strategy yields strong results, exploring more advanced fusion techniques is a promising avenue for future research.
@@
Figure 4: Teacher Fusion across varying maximum ranges from the current sample. Baseline refers using only strong augmentations.
IV-F Combining Components
@@
Building on the above results, we further explore how these components interact in SSL for online mapping. Starting with the teacher-student setup as a Core we incrementally add components and evaluate their impact on performance. Here, we add strong augmentations to the student, +Augs, with Photometric, CutOut, and BEVDrop. Teacher fusion, +Fusion, is used with a maximum range of $30$m. The Cosine feature similarity loss is incorporated under +Featsim with a weight of $w_{\text{feat}}=0.25$. Additionally, a threshold of $0.6$ is used without any temperature sharpening in +Thr, and hard pseudo-labeling is further added and denoted with +Hard.
@@
The results in Tab. III demonstrate that combining components leads to further increased performance. Adding strong augmentations to the student yields a significant performance boost, as does further incorporating teacher fusion. Together these contribute to mIoU gains of $16.1$ and $15.1$ for $2.5\%$ and $10\%$ respectively. When combined with Fusion, feature similarity slightly improves performance for $2.5\%$ labelled data, but has no impact for $10\%$. Finally, applying thresholding yields a modest improvement for both ratios, while using the hard pseudo-labels does not contribute further. These results underscore the importance of conducting ablation studies across different labelled data ratios, as the impact of various modules can vary depending on the amount of labelled data.
TABLE III: mIoU when incrementally adding SSL components.
Label Util. Core +Augs +Fusion +Featsim +Thr. +Hard
@@
IV-G Varying Label Utilisation
We are also interested in how our proposed SSL method scales with varying amounts of labelled data. Using the parameters from Sec. IV-F, our proposed method includes strong augmentations, teacher fusion, Cosine feature similarity, and thresholding. Tab. IV displays performance when using both unlabelled and labelled data against using only the labelled data. Each setup is repeated three times with different data folds. As before, we report the mean and standard deviation; however, in this case, we select the best-performing model on the validation set and evaluate it on the test set. Besides conducting experiments on AV2 [45], we also use NuScenes [3], using the same hyperparameters except for learning rate and weight decay. As expected, performance decreases as the amount of labelled data is reduced. Here, we experiment using $100\%$ of the data, corresponding to $700$ sequences, down to only $18$ sequences at $2.5\%$ utilisation. Using the unlabelled data and our proposed setup yields significant performance gains across all labelled data utilisation ratios. In the lowest setting, relying solely on the labelled sequences results in poor performance, with mIoU dropping as low as $7.5$ for AV2 and $5.6$ for nuScenes. However, incorporating unlabelled data significantly boosts performance to $28.9$ and $15.5$ mIoU for the two datasets respectively. Using only $10\%$ of the labels, the performance gap between using the unlabelled data and not is reduced from $29.6$ mIoU to $3.5$ mIoU on AV2, and from $12$ mIoU to $3.4$ mIoU on NuScenes.

### v2

@@ 3.6 Teacher Fusion
To ensure that additional frames contribute meaningful new information, we randomly sample them at varying distances relative to the current frame. This strategy mitigates issues with strictly time-based selection, such as: ego vehicle being stationary, which provides no new information, and the vehicle moving at high speeds, leading to minimal spatial overlap between frames.
4 Experiments
Table 1: Performance (mIoU) of various augmentations. Using multiple augmentations yields strong results, significantly outperforming no augmentations. We also note that CutOut [12] works better than CamDrop [22] for online mapping.
{NiceTabular}
[…]
@@ 4 Experiments
✓ ✓ $21.9^{\pm 1.0}$ $31.9^{\pm 0.1}$
✓ ✓ $24.4^{\pm 1.2}$ $33.9^{\pm 0.2}$
✓ ✓ ✓ $\textbf{25.5}^{\pm 0.7}$ $\textbf{34.3}^{\pm 0.1}$
Figure 2: No thresholding, soft MixMatch [2], and hard FixMatch[41] thresholding utilising $2.5\%$ and $10\%$ of the labels. Baseline, mean and std., use neither thresholding nor sharpening.
To explore the effectiveness of SSL for online mapping, we describe the experimental setup in Sec. 4.1 and conduct comprehensive ablation studies in Sec. 4.2 to evaluate the contribution of each component. Building on these results, we further evaluate the performance in Sec. 4.3.
4.1 Experimental Setup
We employ a ResNet50 [20] as image encoder $f^{\text{enc}}_{\theta}$, LSS [39] for lifting features to BEV $f^{\text{lift}}_{\theta}$, and a lightweight convolutional bottleneck decoder inspired by [19] as $f^{\text{dec}}_{\theta}$. We perform multi-label prediction for pedestrian crossings, lane dividers, and road boundaries and use Focal Loss [29] as the supervised loss $\ell_{\text{sup}}$ on the labelled dataset and $\ell_{\text{cls}}$ for the unlabelled data. The region of interest follows [31] spanning ±$\pm 45$m longitudinally and ±$\pm 15$m laterally. Each cell has a side length of $0.3$m. For brevity, we report the mean Intersection over Union (mIoU) across these three classes. To account for the stochasticity in the results from individual runs, we either report the mean and standard deviation of the mIoU for three runs with the same hyperparameters or multiple nearby values of the hyperparameter. We choose to follow [15] and use unlabelled weight $\omega_{\text{cls}}=1$. We follow [22] and linearly ramp-up $\omega_{\text{cls}}$, and when applicable $\omega_{\text{feat}}$, over the first third of the training. Unless stated otherwise, no thresholding or sharpening is used in the experiments. For the CutOut augmentation, we remove $25\%$ of the image, and for BEV feature dropout (BEVDrop) we follow [22, 52] and drop $50\%$ of the features. The models, unless stated otherwise, are trained and evaluated using the Argoverse 2 (AV2) dataset [49] with the geographically disjoint splits proposed by [28] using surround-view camera images as input and use the provided HD maps as labels. We generally experiment using two utilisation rates of labelled data, namely $2.5\%$ and $10\%$, corresponding to $18$ and $70$ labelled sequences respectively.
4.2 Exploration
Table 2: mIoU for sharpening methods using threshold $0.6$. Neither soft MixMatch [2] nor hard thresholding FixMatch[41] show significantly impact over using no thresholding
{NiceTabular}
[…]
@@ 4.2 Exploration
✓ ✓ $25.4^{\pm 0.8}$ $34.6^{\pm 0.1}$
We explore different SSL approaches outlined in Sec. 3 to show their effectiveness for segmentation-based online mapping.
Augmentations: We evaluate various strong augmentation strategies for the student model, as discussed in Sec. 3.3. The results in Sec. 4 demonstrate that increasing the difficulty of the task for the student model is crucial also for online mapping. With $10\%$ label utilisation, Photometric augmentation alone is sufficient for significant model improvement. However, with $2.5\%$ label utilisation, additional augmentations are required. For both label utilisation ratios, applying Photometric, CutOut, and BEVDrop yields the best performance, and we adopt this configuration as the Baseline in further experiments. Specifically, we observe a significant performance improvement, with mIoU increasing from $12.1$ to $25.5$ for the $2.5\%$ label utilisation and $21.3$ to $34.3$ for the $10\%$ label utilization. Furthermore, contrary to the findings in [22], we find that removing parts of the images, i.e. CutOut, yields better performance than dropping an entire camera and masking its field of view in the loss, as done with CamDrop.
Sharpening and Thresholding: The effects of sharpening and thresholding are illustrated in Fig. 2. As shown to the left, supervising with the most confident pseudo-labels does not yield significant performance improvements. However, using a low confidence threshold provides a slight boost in performance, and we set the threshold to 0.6 for subsequent experiments. The right plot shows the impact of sharpening. While sharpening can lead to minor improvements, the gains are marginal and introduce an additional hyperparameter, making it less appealing for our setting. Sec. 4.2 presents results across multiple runs, comparing hard and soft label supervision. Overall, these techniques have limited impact on online mapping, contrasting findings in other domains where sharpening has been highly effective—typically when used alongside high thresholds [41].
@@
Feature Consistency: We compare the performance using Mean Squared Error (MSE) and Cosine Similarity (COS) losses with $\omega_{\text{cls}}=1$ and varying feature loss weight $\omega_{\text{feat}}$. In Fig. 3 the results are displayed using features from before or after the BEV decoder $f^{\text{dec}}_{\theta}$ as inputs to $\ell_{\text{feat}}$. When utilizing $2.5\%$ of the labels the results are noisy, and no clear benefit from the additional supervision is observed across the setups. However, for the $10\%$ label utilisation, late COS consistency provides a slight performance gain and demonstrates greater stability than the MSE loss. While prior BEV segmentation works opted to use MSE [60, 15], we believe COS to be the better choice for our setting.
Figure 3: Feature Similarity using Mean Square Error following [60] and Cosine Similarity following [46] varying $\omega_{\text{feat}}$. Baseline refers to no feature similarity while early and late refers to using BEV features before or after BEV processing $f^{\text{dec}}_{\theta}$ in Fig. 1.
@@
Teacher Fusion: To enhance pseudo-labels, we propose to fuse teacher predictions from multiple samples. We use two additional samples, as we empirically found, see Sup. mat. Sec. 6, that adding more samples does not provide further benefits but makes training slower due to the additional inference costs. Therefore, we focus on examining the influence of the maximum distance to these additional samples. The results in Fig. 4 show that even our simple fusion of predicted probabilities generally improves performance, both with and without thresholding. The fusion of BEV features, on the other hand, shows no significant benefit. For $2.5\%$ label utilisation we observe that increasing the range to above $20$m is required to improve performance, whereas for $10\%$ the maximum range is not as important. In further experiments, we will use $30$m as it works well for both utilisation ratios. While the simple probability fusion strategy yields strong results, exploring more advanced fusion techniques is a promising avenue for future research.
@@
Figure 4: Teacher Fusion across varying maximum ranges from the current sample. Fusing the probabilities yields consistent improvements over the Baseline without any fusion.
@@
Combining Components: Building on the above results, we further explore how these components interact in SSL for online mapping. Starting with the teacher-student setup as a Core we incrementally add components and evaluate their impact on performance. Here, we add strong augmentations to the student, +Augs, with Photometric, CutOut, and BEVDrop. Teacher fusion, +Fusion, is used with a maximum range of $30$m. The Cosine feature similarity loss is incorporated under +Featsim with a weight of $w_{\text{feat}}=0.25$. Additionally, a threshold of $0.6$ is used without any temperature sharpening in +Thr, and hard pseudo-labeling is further added and denoted with +Hard.
@@
The results in Tab. 3 demonstrate that combining components leads to further increased performance. Adding strong augmentations to the student yields a significant performance boost, as does further incorporating teacher fusion. Together these contribute to mIoU gains of $16.1$ and $15.1$ for $2.5\%$ and $10\%$ respectively. When combined with Fusion, feature similarity slightly improves performance for $2.5\%$ labelled data, but has no impact for $10\%$. Finally, applying thresholding yields a modest improvement for both ratios, while using the hard pseudo-labels does not contribute further. These results underscore the importance of conducting ablation studies across different labelled data ratios, as the impact of various modules can vary depending on the amount of labelled data.
Table 3: mIoU when incrementally adding SSL components.
Lbl. Util. Core +Augs +Fusion +Featsim +Thr. +Hard
@@
4.3 Putting SSL to the Test
Building on the design choices made above, we aim to further verify the strengths of our proposed SSL scheme. Using the parameters from Fig. 4, our proposed method includes strong augmentations, teacher fusion, cosine feature similarity, and thresholding. As before, we report the mean and standard deviation; however, in this case, we select the best-performing model on the validation set and evaluate it on the test set. The performance is compared by utilizing both unlabelled and labelled data with the SSL-scheme versus using only the labelled data in a supervised setting. First, we compare our SSL scheme against swapping the backbone for a vision foundation model. Next, we evaluate the robustness of our method across different online mapping architectures. We then examine its scalability with varying amounts of labelled data and demonstrate the method’s effectiveness for domain adaptation.
Comparing with pre-training: We substitute the ImageNet pre-trained ResNet50 with DINOv2 [36] vision foundation model, which leverages pre-training on a substantially larger dataset. Although DINOv2’s size makes it impractical for vehicular deployment, it demonstrates performance improvements in Tab. 4. Nevertheless, a significant performance gap persists when comparing DINOv2 without SSL to our SSL approach. Notably, integrating our SSL scheme with the DINOv2 backbone eliminates this gap, underscoring the effectiveness of our approach.
Table 4: mIoU when when using DINOv2 backbone adding SSL.
Backbone SSL Label utilisation
$2.5\%$ $10\%$
ResNet50 $7.5^{\pm 0.9}$ $10.6^{\pm 0.2}$
DINOv2 $11.3^{\pm 0.6}$ $17.3^{\pm 0.4}$
ResNet50 $29.4^{\pm 0.3}$ $36.8^{\pm 0.2}$
DINOv2 $28.7^{\pm 0.9}$ $33.2^{\pm 0.2}$
Robustness to model architecture: We apply our proposed SSL setup to state-of-the-art online mapping models. As shown in Fig. 5, various architectures—including CVT [59], IPM [19], GKT [8], and LSS [39] – all benefit significantly from the SSL scheme. LSS achieves the highest overall performance, with a $3.9$x and $3.5$x boost when $2.5\%$ and $10\%$ of the data have labels, respectively. While e.g. CVT performs slightly worse overall, it sees even greater relative improvements, with performance increasing by $4.8$x and $4.6$x under the same label utilizations.
Figure 5: Multiple SOTA online mapping methods benefit from the SSL-scheme compared to the supervised training.
Scalability:
(a) Argoverse 2
(b) nuScenes
Figure 6: Test performance (mIoU) across utilising various amount of the labelled nuScenes data. All data is used as unlabelled for all runs. Training with unlabelled data using the SSL-scheme consistently improves performance compared to only using supervised training across all label utilisations.
Figure 7: Qualitative results with varying amounts of labelled data. Predictions are shown for models trained with $10\%$, $25\%$, and $100\%$ of the labelled data, using both standard supervised training and our proposed SSL approach.
Figure 8: Adapting to new cities for AV2. Adding unlabelled sequences with the SSL-scheme from the target domain enhances performance in the target city. When adapting to Pittsburgh, the gap is almost completely closed to using all target labels.
We conduct experiments on AV2 [49] and nuScenes [3], with the same hyperparameters except for learning rate and weight decay. Our experiments range from using the labels from the full dataset, corresponding to $700$ sequences, down to just $18$ sequences at $2.5\%$ utilisation. Figure 6 shows how our proposed SSL method and its supervised counterpart scale with varying amounts of labelled data. For the supervised models, performance decreases rapidly as the amount of (labelled) data is reduced. In the most label-scarce setting, having access to $2.5\%$ of the labels, models trained solely on labelled sequences perform poorly, with mIoU dropping to $7.5$ for AV2 and $5.6$ for nuScenes. Incorporating unlabelled data with our proposed approach yields substantial performance boosts across all labelled data utilisation ratios. Leveraging unlabelled data with the SSL-scheme dramatically improves performance to $28.9$ and $15.5$ mIoU for the respective datasets.
If $10\%$ of the data has labels, our method achieves a $3.5$x performance boost compared to only using the labelled data. This narrows the gap to a fully supervised model that uses all labels, to just $3.5$ mIoU on AV2. For nuScenes, the performance boost is $1.5$x, and the remaining gap to when all labels have been used is reduced to $3.5$ mIoU.
To provide visual context for these quantitative improvements, Fig. 7 shows predictions from models trained only on labelled data against those trained with both labelled and unlabelled data. The predictions are noticeably more stable and reliable when using our semi-supervised method, aligning with the quantitative improvements reported above.
City Adaptation: We assess our method’s ability to adapt to new cities using the AV2 Far Extrapolation splits from [28] and the Boston-to-Singapore split from [35] for nuScenes. Our approach utilises all labelled data from the source cities while gradually incorporating unlabelled data from the target cities, ensuring geographic separation from the test set. As shown in Fig. 8 and Fig. 9 for AV2 and nuScenes, respectively, we find that even without labelled data from the target city, performance improves significantly by utilizing its unlabelled data. The domain gap between Boston and Singapore in nuScenes is substantial, making adaptation more challenging than in AV2, where all data comes from US cities. Notably, performance improves as more unlabelled sequences from the target domain are added, suggesting that additional unlabelled data could further enhance results. For smaller domain gaps, such as adapting to Pittsburgh in AV2, our method nearly matches the performance of a fully supervised model, with only a $0.5$ mIoU difference. For larger domain gaps, such as Boston to Singapore, our method boosts mIoU from $11.1$ to $17.0$ when using $100$ unlabelled sequences from Singapore. However, a substantial gap remains in cases of large domain shifts, posing a key challenge for future research.
Figure 9: City adaptation from Boston to Singapore on nuScenes. While performance naturally is lower than using all Singapore labels, incorporating unlabelled sequences using the SSL-scheme helps improve the model’s performance when labelled data is only available from Boston.
5 Conclusion
This paper draws inspiration from SSL techniques in other domains, demonstrating their effectiveness for segmentation-based online mapping. We show that applying photometric, CutOut, and BEVDrop augmentations in a teacher-student setup already performs well. When combined with our proposed multi-step fusion, and feature similarity constraints we show an additional performance boost.
Using the proposed SSL-scheme, we show that incorporating unlabelled data substantially improves performance on both AV2 and nuScenes compared to relying solely on labelled data. For example, if $10\%$ of the data has labels, the SSL-approach to leverage unlabelled data achieves a $3.5$x performance boost compared to only using the labelled data, narrowing the gap to a fully supervised model that uses all labels to just $3.5$ mIoU. We also show strong generalization to unseen cities and believe that additional unlabelled data could further enhance the domain adaptation results. Still, when adapting to Pittsburgh (AV2), incorporating purely unlabelled target-domain data reduces the performance gap from $5$ to just $0.5$ mIoU.
In summary, efficiently leveraging both labelled and unlabelled data is crucial for advancing online mapping. While this work does not explore all existing SSL methods, it provides a foundation for understanding key SSL methods in this context and highlights their potential. We hope this study inspires further research in teacher fusion strategies, self-supervised pre-text tasks, and broader semi-supervised learning approaches to enhance online mapping performance further.
Acknowledgements
We want to express our sincere gratitude to Adam Tonderski, Georg Hess, William Ljungbergh, Carl Lindström, Ji Lan, and Erik Stenborg for their continuous support and invaluable discussions. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Berzelius and C3SE Alvis partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Supplementary Material
6 Exact numbers
For completeness, we also show the exact performance numbers for the results shown in Figures in Sec. 4. For thresholds and temperatures, the numbers are displayed in Tab. 5 and Tab. 6 for $2.5\%$ of the labels, respectively. For $10\%$, we refer to Tab. 7 and Tab. 8. In addition to these, the performance for feature similarity loss with different weights is shown in Tab. 9 and Tab. 10 for $2.5\%$ and $10\%$ label utilisation, respectively. The validation performance for teacher fusion over various ranges is provided in Tab. 11 and Tab. 12 for $2.5\%$ and $10\%$ label utilisation. We also experiment with the amount of samples to use in addition to the center sample in Tab. 13. For the test performance across multiple label utilisations, the numbers for both Argoverse 2 (AV2) and nuScenes are shown in Tab. 14. Finally, Tab. 15 provides the performance for city adaptation for both Argoverse 2 and nuScenes when varying the amount of unlabelled data.
Table 5: Validation performance (mIoU) for different thresholds on Argoverse 2 $2.5\%$ of the labels utilised.
Thresholds $0.55$ $0.6$ $0.65$ $0.7$ $0.75$ $0.8$ $0.85$ $0.9$
Thr. $24.1$ $26.8$ $25.9$ $23.6$ $25.7$ $23.6$ $24.6$ $22.3$
Thr.+Hard $24.4$ $26.6$ $25.2$ $25.5$ $24.2$ $25.0$ $23.7$ $24.6$
Table 6: Validation performance (mIoU) for using sharpening with different temperatures on Argoverse 2 $2.5\%$ of the labels utilised.
Temperature $0.05$ $0.1$ $0.25$ $0.5$ $0.75$ $0.95$
No Thr. $27.6$ $25.3$ $27.4$ $27.1$ $27.0$ $26.1$
Thr. $24.3$ $26.2$ $27.4$ $25.3$ $25.7$ $26.2$
Thr.+Hard $24.2$ $25.1$ $25.3$ $26.3$ $26.9$ $24.4$
Table 7: Validation performance (mIoU) for different thresholds on Argoverse 2 $10\%$ of the labels utilised.
Thresholds $0.55$ $0.6$ $0.65$ $0.7$ $0.75$ $0.8$ $0.85$ $0.9$
Thr. $33.6$ $33.9$ $33.4$ $33.3$ $33.0$ $32.4$ $32.1$ $30.9$
Thr.+Hard $34.8$ $34.7$ $34.4$ $33.7$ $33.3$ $32.3$ $31.4$ $31.4$
Table 8: Validation performance (mIoU) for using sharpening with different temperatures on Argoverse 2 $10\%$ of the labels utilised.
Temperature $0.05$ $0.1$ $0.25$ $0.5$ $0.75$ $0.95$
No Thr. $35.0$ $34.8$ $34.9$ $34.7$ $33.8$ $34.6$
Thr. $34.2$ $35.3$ $34.8$ $34.2$ $32.9$ $33.5$
Thr.+Hard $34.6$ $34.8$ $34.4$ $34.9$ $33.9$ $34.1$
Table 9: Validation performance (mIoU) for feature similarity loss with different weights on Argoverse 2 $2.5\%$ of the labels utilised.
Weight $0.05$ $0.1$ $0.25$ $0.5$ $1.0$ $1.5$ $2.0$ $2.5$
MSE-pre $24.3$ $25.4$ $26.0$ $22.2$ $23.3$ $20.0$ $23.1$ $22.5$
MSE-post $24.3$ $22.5$ $22.5$ $23.1$ $24.3$ $22.5$ $23.4$ $24.1$
COS-pre $25.3$ $25.3$ $24.8$ $25.0$ $23.4$ $24.9$ $24.6$ $22.9$
COS-post $23.4$ $25.1$ $23.4$ $23.6$ $25.5$ $26.1$ $24.4$ $25.4$
Table 10: Validation performance (mIoU) for feature similarity loss with different weights on Argoverse 2 $10\%$ of the labels utilised.
Weight $0.05$ $0.1$ $0.25$ $0.5$ $1.0$ $1.5$ $2.0$ $2.5$
MSE-pre $33.7$ $32.8$ $32.3$ $32.5$ $32.4$ $30.0$ $28.8$ $24.1$
MSE-post $33.6$ $33.4$ $32.3$ $31.9$ $31.2$ $29.3$ $28.0$ $28.0$
COS-pre $34.4$ $33.8$ $34.4$ $34.4$ $34.1$ $33.8$ $34.4$ $33.2$
COS-post $34.0$ $34.8$ $35.4$ $35.1$ $34.4$ $34.8$ $33.4$ $31.4$
Table 11: Validation performance (mIoU) for teacher fusion over various ranges on Argoverse 2 $2.5\%$ of the labels utilised.
Ranges $10$ $15$ $20$ $25$ $30$ $35$ $40$
Probs. $27.7$ $27.7$ $27.2$ $27.6$ $28.4$ $28.7$ $26.8$
Probs.+Thr. $25.5$ $25.9$ $27.3$ $28.2$ $26.9$ $27.5$ $26.7$
Feats. $24.1$ $25.6$ $24.8$ $23.6$ $23.1$ $25.8$ $26.2$
Feats.+Thr. $26.5$ $26.6$ $27.0$ $26.5$ $26.1$ $26.7$ $25.6$
Table 12: Validation performance (mIoU) for teacher fusion over various ranges on Argoverse 2 $10\%$ of the labels utilised.
Ranges $10$ $15$ $20$ $25$ $30$ $35$ $40$
Probs. $36.2$ $35.5$ $36.5$ $36.1$ $35.5$ $35.8$ $35.9$
Probs.+Thr. $35.9$ $36.2$ $36.4$ $35.4$ $35.7$ $35.8$ $36.0$
Feats. $34.0$ $34.5$ $34.2$ $34.4$ $34.3$ $34.0$ $34.4$
Feats.+Thr. $34.2$ $34.0$ $34.8$ $34.1$ $34.4$ $35.4$ $34.6$
Table 13: Multi-step Fusion with varying numbers of additional samples within a 20m range using $10\%$ of labels. The best results are achieved with 4 additional samples.
Additional samples
$2$ $4$ $6$
Probs $34.6$ $35.0$ $30.2$
Feats $32.4$ $32.8$ $26.1$